OpenStudy (anonymous):

How would you work out this question? A footballer scored 15 goals over 6 games. How many games will it take for him to have scored 50 goals?

OpenStudy (ghazi):

simply 15 goals in 6 games 1 goal in = 6/15 games and 50 goals= 6/15*50 games= 20 games

OpenStudy (ghazi):

is it clear?

OpenStudy (anonymous):

yes thank you very much